Identify the rate in the statement " 260 is 52% of 500."
A) 500 B) 260 C) 52%

Answer by josmiceli(6793)

c) 52% is the answer
Arate is never just a number. It is a ratio that can look different
depending on the numbers you use for the ratio.
52% = 52/100 = 26/50 = 13/25 = 130/250 = .52
each one is a ratio. Note that 52% is really 52/100